OpenStudy (anonymous):

Do you mind helping me out with this one as well @SavannahMiller27 Can you please solve the inequality below and graph the solution on a number line –2(x – 3) ≥ 5 – (x + 3) Do you mind helping me out with this one as well @SavannahMiller27

OpenStudy (anonymous):

Yeah I'll help you.

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

Okay 1st thing you would do is distribute the -2 to the (x-3) Written out it looks like this: \[(-2 \times x ) ( -2 \times -3 )\] so it would be: \[-2x + 6 \ge 5 - (x+3)\]

OpenStudy (anonymous):

The next thing I would do is distribute the negative to the (x + 3). so then it would be (-x - 3)

OpenStudy (anonymous):

Next you would simplify both sides. You can simplify the right side to: -x + 2

OpenStudy (anonymous):

So your inequality now looks like this: \[-2x + 6 \ge -x + 2\]

OpenStudy (anonymous):

Then what I would do is add x to both sides. \[-2x + x + 6 \ge -x +2 + x\] which would simplify to this: \[-x + 6 \ge 2\]

OpenStudy (anonymous):

So now what you would do is subtract 6 from both sides. \[-x + 6 -6 \ge 2 -6\] simplify: \[-x \ge -4 \]

OpenStudy (anonymous):

x cannot be negative so divide by -1. Remember when you divide by a negative number you have to switch the sign. \[-x \div -1 \le -4 \div-1\] simplify to: \[x \le 4 \]

OpenStudy (anonymous):